OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 - 13x + 42 A. x2 + 12 B. x + 84 C. x + 7 D. x - 7

OpenStudy (anonymous):

(x-7)(x-6) is what it factors to, so D

OpenStudy (anonymous):

thanks :)

OpenStudy (jackellyn):

x^2 - 13x + 32 = 0 (x - 6)(x - 7) = 0. x = 6 AND x = 7.